# Bridging Constraints and Stochasticity: A Fully First-Order Method for Stochastic Bilevel Optimization with Linear Constraints

### v1

@@ 2 Related Work
Nonsmooth and nonconvex optimization. The nonsmooth and nonconvex structure of bilevel optimization with constraints makes its analysis closely related to nonsmooth nonconvex optimization. The best-known convergence result is given by zhang2020complexity, which establishes optimal convergence rates of $O(\delta^{-1}\epsilon^{-3})$ for the deterministic case and $O(\delta^{-1}\epsilon^{-4})$ for the stochastic case. Our result of $\tilde{O}(\delta^{-1}\epsilon^{-5})$ is one factor of $\epsilon$ away from the optimal stochastic rate, indicating potential room for improvement.
@@ 4.1 Stochastic Implementation
We compute the stochastic hypergradient oracle via a penalty formulation with smooth activation as
@@
3: Compute $\tilde{y}^{*}(x)$ and $\tilde{\lambda}(x)$ by SGD such that $\|\tilde{y}^{*}(x)-y^{*}(x)\|\leq O(\delta)$ and $\|\tilde{\lambda}(x)-\lambda^{*}(x)\|\leq O(\delta)$
4: Define the smooth Lagrangian $L_{\tilde{\lambda},\alpha}(x,y)$ using the penalty Lagrangian and smooth activation function $\rho(x)$
5: Compute $\tilde{y}(x)=\operatorname*{arg\,min}_{y}L_{\tilde{\lambda},\alpha}(x,y)\|\leq\delta$ by SGD
@@ Appendix A Appendix
$\|\nabla h\|\leq\|\mathbf{A}\|+\|\mathbf{B}\|\leq 2M_{AB}$ (13)
@@
$\|\alpha_{1}(\nabla h)^{T}\Delta\lambda\|\leq\alpha_{1}\cdot 2M_{AB}\cdot C_{\lambda}\delta=O(\alpha_{1}\delta)$ (14)
@@
$\implies|h_{i}(x,y)-h_{i}(x,\tilde{y}^{*}(x))|\leq M_{AB}\delta$ (17)
Using (17): $|h_{i}(x,y)|\leq O(\delta)+M_{AB}\delta=O(\delta)$ for near-active constraints.
With $|\Delta\rho_{i}|\leq 1$ and $\|\nabla h_{i}\|\leq 2M_{AB}$:
@@
oracle calls give $\|\,\tilde{y}^{*}-y^{*}\|,\;\|\,\tilde{\lambda}^{*}-\lambda^{*}\|\leq\delta$.
@@
[…]
@@
Since $M=\lfloor\frac{\delta}{D}\rfloor$, we have $M\approx\frac{\delta}{D}$, which gives us:
@@
Let’s set $D=\Theta(\delta\epsilon^{2})$ and $M=\Theta\left(\frac{1}{\epsilon^{2}}\right)$ to satisfy this constraint. From (86), we can determine $K$:

### v2

@@ 2 Related Work
Nonsmooth and nonconvex optimization.
The nonsmooth and nonconvex structure of bilevel optimization with constraints makes its analysis closely related to nonsmooth nonconvex optimization. The best-known convergence result is given by zhang2020complexity, which establishes optimal convergence rates of $O(\delta^{-1}\epsilon^{-3})$ for the deterministic case and $O(\delta^{-1}\epsilon^{-4})$ for the stochastic case. Our result of $\tilde{O}(\delta^{-1}\epsilon^{-5})$ is one factor of $\epsilon$ away from the optimal stochastic rate, indicating potential room for improvement. Note that $\delta$ and $\epsilon$ are independent parameters: $\delta$ controls the radius of the Goldstein subdifferential approximation, while $\epsilon$ controls the stationarity accuracy. In our analysis, we treat $\delta$ as a fixed parameter (not scaling with $\epsilon$), which is consistent with the standard treatment in nonsmooth optimization literature.
@@ 3.1 Assumptions
    Note: The combination of strong convexity and bounded gradients is standard in bilevel optimization. Strong convexity ensures uniqueness of the lower-level solution, while bounded gradients are naturally satisfied when the domain $X\times Y$ is compact or when $g$ has bounded sublevel sets. This assumption is consistent with prior work on constrained bilevel optimization kornowski2024.
@@ 4.1 Stochastic Implementation
We compute the stochastic hypergradient oracle via a penalty formulation with smooth activation. The oracle takes a point $x$ and accuracy parameter $\alpha$, and returns a stochastic approximation $\nabla\tilde{F}(x)$ of the true hypergradient $\nabla F(x)$. The key steps are: (1) approximately solve the constrained lower-level problem to obtain $\tilde{y}^{*}(x)$ and dual variables $\tilde{\lambda}(x)$; (2) construct a smoothed penalty Lagrangian $L_{\tilde{\lambda},\alpha}(x,y)$; (3) minimize this Lagrangian to get $\tilde{y}(x)$; and (4) estimate the hypergradient by averaging stochastic gradient samples.
@@
3: Compute $(\tilde{y}^{*}(x),\tilde{\lambda}(x))$ using SGD (Lemma 4.5) s.t. $\mathbb{E}[\|\tilde{y}^{*}(x)-y^{*}(x)\|]\leq O(\delta)$ and $\mathbb{E}[\|\tilde{\lambda}(x)-\lambda^{*}(x)\|]\leq O(\delta)$
4: Define the smooth Lagrangian $L_{\tilde{\lambda},\alpha}(x,y)$ from Eq 4 and smooth activation function $\rho(x)$
5: Compute $\tilde{y}(x)=\operatorname*{arg\,min}_{y}L_{\tilde{\lambda},\alpha}(x,y)$ by SGD such that $\|\tilde{y}(x)-y^{*}_{\tilde{\lambda},\alpha}(x)\|\leq\delta$
@@ Appendix A Appendix
$\|\nabla h\|\leq\|\mathbf{A}\|+\|\mathbf{B}\|\leq 2M_{\nabla h}$ (13)
@@
$\|\alpha_{1}(\nabla h)^{T}\Delta\lambda\|\leq\alpha_{1}\cdot 2M_{\nabla h}\cdot C_{\lambda}\delta=O(\alpha_{1}\delta)$ (14)
@@
$\implies|h_{i}(x,y)-h_{i}(x,\tilde{y}^{*}(x))|\leq M_{\nabla h}\delta$ (17)
Using (17): $|h_{i}(x,y)|\leq O(\delta)+M_{\nabla h}\delta=O(\delta)$ for near-active constraints.
With $|\Delta\rho_{i}|\leq 1$ and $\|\nabla h_{i}\|\leq 2M_{\nabla h}$:
@@
$y_{t+1}$ $=y_{t}-\eta_{y}\left(\nabla_{y}\tilde{g}(x,y_{t};\zeta_{t})+\mathbf{B}^{T}\lambda_{t}\right),$ (45)
$\lambda_{t+1}$ $=\max\left\{0,\lambda_{t}+\eta_{\lambda}\left(\mathbf{A}x-\mathbf{B}y_{t+1}-\mathbf{b}\right)\right\},$ (46)
where $\eta_{y},\eta_{\lambda}>0$ are step sizes, and $\tilde{g}$ denotes the stochastic gradient oracle. Under Assumption 3.3, this SPD algorithm with appropriate step sizes satisfies linear convergence in expectation: $\mathbb{E}\|y_{t}-y^{*}\|^{2}\leq\bigl(1-\tfrac{1}{\kappa_{g}}\bigr)^{t}\!D_{0}^{2},\quad\kappa_{g}:=C_{g}/\mu_{g}$ and similarly for the dual variables. Hence
@@
oracle calls give $\mathbb{E}[\|\,\tilde{y}^{*}-y^{*}\|],\;\mathbb{E}[\|\,\tilde{\lambda}^{*}-\lambda^{*}\|]\leq\delta$.
@@
A linear-rate variance-reduced method (SVRG) requires $t_{2}=O\!\bigl(\kappa_{\mathrm{pen}}\log(1/\delta)\bigr)=O\!\bigl(\alpha^{-2}\log(1/\delta)/\mu_{g}\bigr)$ oracle calls to attain $\|\tilde{y}-y^{*}_{\tilde{\lambda}^{*},\alpha}\|\leq\delta$(PMLR-SVRG2013).
@@
Since $M=\lfloor\frac{\delta}{D}\rfloor$, we require $M\approx\frac{\delta}{D}$ to be consistent. Combining this with $M=\Theta(1/\epsilon^{2})$ from (89), we obtain:
@@
Note that $\delta$ and $\epsilon$ are independent parameters: $\delta$ controls the radius of the Goldstein subdifferential (a fixed parameter of the stationarity concept), while $\epsilon$ controls the accuracy of stationarity. The relationship $D=\Theta(\delta\epsilon^{2})$ ensures that the algorithm’s clipping parameter $D$ scales appropriately with both parameters. Set $D=\Theta(\delta\epsilon^{2})$ and $M=\Theta\left(\frac{1}{\epsilon^{2}}\right)$ to satisfy this constraint.
From (88), we can determine $K$: